OpenStudy (anonymous):

Solve the following system of equations by using the elimination method. -4x – 2y = -12 4x + 8y = -24

OpenStudy (anonymous):

add both equation. tell me what you got.

OpenStudy (anonymous):

@breeskee03

OpenStudy (anonymous):

6y=-36

OpenStudy (anonymous):

Solve from there.

OpenStudy (anonymous):

-4x – 2y = -12 4x + 8y = -24 ------------------ + 6y = -36 y = ?? input y into 4x+8( ??)=-12 x= ?

OpenStudy (anonymous):

ok by adding both equations you get. 6y=-36 or y=-6 put this value of y in any of the given equations and you will get x. i hope it is clear.

OpenStudy (anonymous):

where is she ? -_-